Hello,
Please help!
Solve the system of equation by elimination.
6x + 5y = -20
-6x - 10y = 25
Thanks

Answer by stanbon(75887)

Solve the system of equation by elimination.
6x + 5y = -20
-6x - 10y = 25
===============
Add to get:
-5y = 5
y = -1
----------
Substitute to solve for "x":
6x + 5*-1 = -20
6x =-15
x = -5/2
===========
Cheers,
Stan H.